ltx2rtf−ConvertsLaTeXfiletoRTFtext-format SYNTAX ltx2rtf[-V][-D][-m](-M][-ooutputfile][-Ccodepage][inputfile] DESCRIPTION Theltx2rtfcommandconvertsaLaTeXfileintoRTFtextformat.Thetextandmuchoftheformatting informationistranslatedtoRTFmakingthenewfilelookverysimilartotheoriginal.If inputfile isnotspecified,standardinputisread.Unlessan outputfile isspecifiedtheresultingRTFisproducedinafileofnametakenfromthesourcewiththe.rtfsuffix. The−Voptionprintsversioninformationonstandardoutputandexits. The−Doptionoutputsverbosedebuggingmessages. The−moptionLaTeX-esequationsandinsertsthemaspictures. The−MoptionLaTeX-esequationsand$embeddedmathsandinsertsthemaspictures. The−C850optionenablesreading850-codedaccentedletters. The−CISOoptionenablesreadingIsoLatin1-codedaccentedletters. UserdefinedLaTexcommandsareignored.IfyouarefamiliarwiththeRTFformatandthecommands canbetranslatedbysimplyinsertingaformatstringintotheRTFfileyoucanaddthesecommandsto thedirect.cfgconfigurationfile.Seeformatdescriptionbelow. UserdefinedLaTeXcommandscanbeexpandedusingexp-macrwhichisprovidedwiththedistribution. config.cfgindicatesthename/locationofprocedurestoLaTeXequationsinthe%TMPDIR%directory andtheconverttheDVIintoaPBMfile. Tocorrectlyconvertfontnamesyoumusteditthefonts.cfgconfigurationfile.Youhavetospecifythefont namesyouuseandhowtheLaTeXdefaultfontnamesshouldbeconvertedtoRTF.Seeformatdescription below. LaTeXvariablesanduserdefinedcommandsarenotevaluatedbyltx2rtfitself.Theywillbesimply ignored.Toletltx2rtfknowthenamesofvariablesyoucanaddthemintheignore.cfgfile.Seeformat descriptionbelow. TheenvironmentvariableRTFPATHmaycontainasearchpathforthesupportfiles(allfilesendingin .cfg).SupportfilesaresearchedfirstintheIfnofileisfoundduringthecurrentdirectory,theninRTF-PATH,theninthedefaultdirectoryspecifiedintheMakefile.Ifthefilesarenotfoundatalltheprogram aborts. \label,\ref,and\pagerefcommandsfromLaTeXaretranslatedtotheRTFfile,butWord6.0badlyhandles\refand\pageref.Thisrequirestocorrectlyedittheresultingfileinthetargetwordprocessor. CAUTION PleasemakesurethattheinputfileisacorrectLaTeXfile.UseLaTeXtofindandcorrecterrorsbefore usingltx2rtf.Theconversioncommanddoesnotproperlyhandleallerrorsbecauseofcorruptinputfiles. Makesurethattheconfigurationfiles direct.cfg,fonts.cfg,config.cfg areinthecorrectdirectory. direct.cfgisusedforconvertingLaTeXcommandsbysimpletextexchange.WritetheLaTexcommand withbackslash,followedbycomma.Therestofthelineuntilthe’.’characterwillbewrittentotheRTF filewhenthecommandisrecognizedintheLaTeXfile.Afterthe’#’charactereverythingisignoreduntil endofline.Toselectaspecificfontuse*fontname*. Makesurethatthefontname fontnameislistedinthefontsconfigurationfileatleastasdummy.Towritethe’*’characteruse"**". example: \alpha,{\f*Symbol*a}.#alphaunderWindowsSymbolFont\copyright,{\ansi\’a9\pc}. fonts.cfgcontainsthefontsconversiontable.AlineconsistsofafontnameinLaTeXfollowedbycomma andafontnameinRTF.Theendismarkedbyadot.Nospacesareallowed.TheLaTeXfontwillbe convertedtotheRTFfontifencounteredintheLaTeXfile.IfmultipletranslationsforaLateXfontare specified,onlythefirstisused.AllfontsinaLaTeXfilethatarenotinthisfilewillbemappedtothe defaultfont.AllRTFfontslistedinthisfilewillbeineveryRTFfileheaderwhetherusedornot.After the’#’charactereverythingisignoredtillendofline.ToaddaRTFfontnotusedassubstitutefora LaTeXfont−forexampleaSymbolfontusedin direct.cfg−useadummyLaTeXnamelikeinthefollowingexample: Dummy3,MathematicalSymbols. Makesureyouusethecorrectfontname.Takecareofspacesinfontnames.Thedefaultfontsarenamed Roman(command\rm),SansSerif(command\sf),Typewriter(command\tt). ignore.cfgisthefilefordefininghowtoignorecertaincommands.Thisfileisusedforrecognitionof LaTeX-variables,userdefinedvariables,andsomesimplecommands.Allvariablesareignoredbutthe convertermustknowthenamestocorrectlyignoreassignmentstovariables.Writethevariable-namewith backslash,followedbycommathenwritethetypeofthevariablefollowedby’.’.PossibleTypesare: NUMBER:simplenumericvalue MEASURE:numericvaluewithfollowingunitofmeasure OTHER:ignoresanythingtothefirstcharacterafter’=’andfromtheretonextspace.eg.\setbox\bak=\hbox COMMAND:ignoresanythingtonext’\’andfromtheretooccurenceofanythingbutaletter.eg.\newbox\bak SINGLE:ignoresinglecommand.eg.\noindent Pleasewritethetypeinuppercaseexactlyasabove.Donotusespaces.Afterthe’#’charactereverythingisignoredtillendofline. Example: \pagelenght,MEASURE. Someenvironmentslikefigurearecurrentlyignored.Itwouldbenicetoscantheincludedinformationand convertthingsifpossible. Theconfigurationfilesarereopenedforeachtoken,thereshouldreallybeaninternalrepresentationofthe configurationfilesthatisreadonceandthenconsultedbymoreefficientmechanisms. SEEALSOrtf2LaTeX(1),LaTeXUser’sGuide&ReferenceManualbyLeslieLamport